Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Input Mode Cleaner cavity)
derossi, spinicelli, gosselin, melo - 17:22 Monday 07 September 2020 (49366) Print this report
IMC Aa sensing matrix measurement

Today we worked on the IMC alignment.

The drift control was not working because the offset on the IB ty error signal was too high, so we aligned at first without the drift control.

We manually tuned the Fmoderr (1st plot), then we moved the IB and the MC to find a good working point for the IMC. Once we found it we manually brought to 0 the offsets in Acl and we closed it.

We then measured the Aa sensing matrix exciting with a 2 Hz line all the DOFs.

DOF Ampl GPS start dur [s]
IB tx 0.8 1283524010 60
IB ty 0.8 1283524100 -
IB tz 0.8 1283524230 -
MC tx 0.1 1283524510 -
MC ty 0.1 1283524613 -

With the new sensing matrix we could close the IMC Aa in full bandwidth (see 2nd plot).

We finally closed the RFC and the PSTAB.

Images attached to this report
Search Help